OpenStudy (anonymous):

Determine whether the function f is even, odd, or neither. x^-6

OpenStudy (anonymous):

Usually if the exponent of a simple power function is even, the function is even, which means the graph looks the same on both sides of the y-axis

OpenStudy (calculusfunctions):

f(x) is even if f(-x) = f(x) f(x) is odd if f(-x) = -f(x) Hence if \[f(x)=\frac{ 1 }{ x ^{6} }\] then first find f(-x). Thus \[f(-x)=\frac{ 1 }{ (-x)^{6} }\] \[=\frac{ 1 }{ x ^{6} }\] Thus f(-x) = f(x) ∴ f(x) is an even function. Do you understand?